OpenStudy (shaunicemorrison):

Given the equation 2Square root of x minus 5 = 2, solve for x and identify if it is an extraneous solution

OpenStudy (er.mohd.amir):

\[2\sqrt{x-5}=2\]

OpenStudy (er.mohd.amir):

eq is this correct

OpenStudy (er.mohd.amir):

\[\sqrt{x-5}=2/2\]

OpenStudy (er.mohd.amir):

\[\sqrt{x-5}=1\]

OpenStudy (er.mohd.amir):

squiring both sides

OpenStudy (er.mohd.amir):

x-5=1 x=1+5 x=6